# Multiple-policy Evaluation via Density Estimation

### v1

@@ 3.1 Contributions
In this work, we propose an algorithm named $\mathrm{CAESAR}$ for multiple-policy evaluation with two phases. In the first phase, we roughly estimate the visitation distributions of target polices with a lower-order sample complexity $\tilde{O}(\frac{1}{\epsilon})$ such that $|\hat{d}^{\pi^{k}}_{h}(s,a)-d^{\pi^{k}}_{h}(s,a)|\leq\max\{\epsilon,\frac{d^{\pi^{k}}_{h}(s,a)}{4}\},k\in[K]$ where $d^{\pi^{k}}_{h}$ is the true visitation distribution of policy $\pi$ at step $h$ and $\hat{d}^{\pi^{k}}_{h}$ is our coarse distribution estimator. In the second phase, with the coarse distribution estimators, we can solve a convex optimization problem to get our optimal sampling distribution $\tilde{\mu}$ with which we estimate the performance of target policies using marginal importance weighting. $\mathrm{CAESAR}$ finally achieves that with number of trajectories $n=\tilde{O}\left(\frac{H^{4}}{\epsilon^{2}}\sum_{h=1}^{H}\max_{k\in[K]}\sum_{s,a}\frac{(d_{h}^{\pi^{k}}(s,a))^{2}}{\mu^{*}_{h}(s,a)}\right)$ and probability at least $1-\delta$, we can evaluate the performance of all target policies up to $\epsilon$ error. $\mathrm{CAESAR}$ is consistently better than the naive uniform sampling strategy over target policies. $\mathrm{CAESAR}$ also improves the results (1) in Dann et al. (2023) in some cases where their results have a dependency on $K$ while we do not (see Section 4.5).
In addition to our main result, we also provide some results that may spark interest beyond the specific multi-policy evaluation problem we addressed in this work. To utilize marginal importance weighting, we propose an algorithm named $\mathrm{IDES}$ to estimate the marginal importance ratio by minimizing a carefully designed step-wise loss function using stochastic gradient descent which is modified from the idea of DualDICE Nachum et al. (2019). We also utilize a Median-of-Means estimator Minsker (2023) to convert the in-expectation result to the high-probability result which can be of interest.
@@ 4 Main Results and Algorithm
In this section, we introduce our algorithm and present our main results. Different from on-policy evaluation, we try to build a single sampling distribution with which we can estimate the performance of all target policies using importance weighting. We achieve it by the following procedures. We first roughly estimate the visitation distributions of target policies at the cost of a lower-order sample complexity. Based on these coarse distribution estimators, we can calculate an optimal sampling distribution by solving a convex optimization problem. Finally, we utilize the idea of DualDICE Nachum et al. (2019) with some modifications to estimate the importance-weighting ratio. We provide the main algorithm scheme in Algorithm 2. In the following sections, we explain these procedures of our algorithm in details.
4.1 Estimation of visitation distributions of target policies
We first introduce a proposition that shows we can roughly estimate the visitation distributions of target policies with just lower-order sample complexity $\tilde{O}(\frac{1}{\epsilon})$. Though, the estimator is coarse and cannot be used to estimate the performance directly which is our ultimate goal, it has some nice properties which enable us to build the optimal sampling distribution and estimate the importance weighting ratio in the following sections.
The idea behind this estimator is based on the following Lemma that shows estimating the mean value of a Bernoulli random variable up to constant multiplicative accuracy only requires $\tilde{O}(\frac{1}{\epsilon})$ samples.
@@ 4.1 Estimation of visitation distributions of target policies
Let $Z_{\ell}$ be i.i.d. samples $Z_{\ell}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathrm{Ber}(p)$, for some known constant $C>0$, setting $t\geq\frac{C\log(C/\epsilon\delta)}{\epsilon}$, we have that with probability at least $1-\delta$, the empirical mean estimator $\hat{p}_{t}=\frac{1}{t}\sum_{\ell=1}^{t}Z_{\ell}$ satisfies,
$|\hat{p}_{t}-p|\leq\max\{\epsilon,\frac{p}{4}\}.$
Proof.
See Appendix A.1. ∎
@@
Proof.
Based on Lemma 4.1, the proposition can be directly derived by regarding $d^{\pi^{k}}_{h}(s,a)$ as a $\mathrm{Bernoulli}$ variable. ∎
@@
Proof.
Proof of the first claim: If $\max\{\epsilon^{\prime},\frac{d(s,a)}{4}\}=\epsilon^{\prime}$, then we have $d(s,a)\leq 4\epsilon^{\prime}$ and
$\hat{d}(s,a)-d(s,a)\leq\epsilon^{\prime},$
hence, $\hat{d}(s,a)\leq d(s,a)+\epsilon^{\prime}\leq 5\epsilon^{\prime}$ which implies that $\hat{d}(s,a)\leq 5\epsilon^{\prime}$ is a necessary condition for $\max\{\epsilon^{\prime},\frac{d(s,a)}{4}\}=\epsilon^{\prime}$.
Proof of the second claim: We have $\hat{d}(s,a)\leq 5\epsilon^{\prime}$. If $\max\{\epsilon^{\prime},\frac{d(s,a)}{4}\}=\epsilon^{\prime}$, then $d(s,a)\leq 4\epsilon^{\prime}$. If $\max\{\epsilon^{\prime},\frac{d(s,a)}{4}\}=\frac{d(s,a)}{4}$, then $d(s,a)\leq\frac{4}{3}\hat{d}(s,a)\leq\frac{20}{3}\epsilon^{\prime}\leq 7\epsilon^{\prime}$. ∎
Based on Lemma 4.3, we can ignore the state-action pairs such that $\hat{d}(s,a)\leq 5\epsilon^{\prime}$. Since if we replace $\epsilon^{\prime}$ by $\frac{\epsilon}{14SA}$, the error of performance estimation induced by ignoring these state-action pair is at most $\frac{\epsilon}{2}$. For simplicity of presentation, we can set $\hat{d}^{\pi}(s,a)=d^{\pi}(s,a)=0$ if $\hat{d}^{\pi}(s,a)<\frac{5\epsilon}{14SA}$. Hence, we have that,
@@ 4.2 Optimal sampling distribution
We evaluate the expected total rewards of target policies by importance weighting using samples $\{s_{1}^{i},a_{1}^{i},s_{2}^{i},a_{2}^{i},\dots,s^{i}_{H},a^{i}_{H}\}_{i=1}^{n}$ sampled from a sampling distribution $\{\mu_{h}\}_{h=1}^{H}$.
From the perspective of minimizing the variance of our estimator (see Appendix A.2), we can find the optimal sampling distribution by solving the following optimization problem,
@@
The above problem is a convex optimization problem. However, the optimal $\mu^{*}_{h}$ can be outside of the convex hull of $\{d^{\pi^{k}}_{h}\}_{k=1}^{K}$. In some cases, the optimal $\mu^{*}$ may not be realized by any policy (see Appendix A.3). Hence, in order to have an easy way of sampling from $\mu^{*}$, we constrain $\mu$ to be inside of the convex hull of $\{d^{\pi^{k}}_{h}\}_{k=1}^{K}$ and solve the constrained optimization problem,
$\mu^{*}_{h}=\operatorname*{arg\,min}_{\mu\in M_{h}}\max_{k\in[K]}\sum_{s,a}\frac{(d_{h}^{\pi^{k}}(s,a))^{2}}{\mu(s,a)},\ h=1,\dots,H,$ (5)
where $M_{h}=\{\mu:\mu=\sum_{k=1}^{K}\alpha_{k}\hat{d}^{\pi^{k}}_{h},\alpha_{k}\geq 0,\sum_{k=1}^{K}\alpha_{k}=1\}$. Denote by $\mu^{*}_{h}=\sum_{k=1}^{K}\alpha^{*}_{k}d^{\pi^{k}}_{h}$ the optimal solution.
Notice that we do not know $d^{\pi}_{h}$, so we can only solve the approximate optimization problem,
$\hat{\mu}^{*}_{h}=\operatorname*{arg\,min}_{\mu\in M_{h}}\max_{k\in[K]}\sum_{s,a}\frac{(\hat{d}_{h}^{\pi^{k}}(s,a))^{2}}{\mu(s,a)},\ h=1,\dots,H,$ (6)
and we denote the optimal solution by $\hat{\mu}^{*}_{h}=\sum_{k=1}^{K}\hat{\alpha}^{*}_{k}\hat{d}^{\pi^{k}}_{h}$. Correspondingly, our real sampling distribution would be $\tilde{\mu}^{*}_{h}=\sum_{k=1}^{K}\hat{\alpha}^{*}_{k}d^{\pi^{k}}_{h}$.
Remark 4.1.
We can constrain $\mu$ to be the convex combination of the visitation distributions of all deterministic policies instead of the target policies. In that case, we can potentially get a better sampling distribution at the cost of exponential dependency of $S$ and $A$ in the lower order sample complexity since we need to roughly estimate the visitation distributions of all deterministic policies. However, we believe there exists a clever way of coarse estimation to get rid of the exponential dependency in the lower order sample complexity.
@@
If property (3) holds: $|\hat{d}_{h}^{\pi^{k}}(s,a)-d^{\pi^{k}}_{h}(s,a)|\leq\frac{d_{h}^{\pi^{k}}(s,a)}{4},\ \forall s\in\mathcal{S},a\in\mathcal{A},h\in[H],k\in[K]$, then
  1. 1.
    $|\tilde{\mu}_{h}^{*}(s,a)-\hat{\mu}_{h}^{*}(s,a)|\leq\frac{\tilde{\mu}_{h}^{*}(s,a)}{4}$ and therefore $\frac{1}{\tilde{\mu}_{h}^{*}(s,a)}\leq\frac{5}{4\hat{\mu}_{h}^{*}(s,a)}$.
  2. 2.
    $\frac{d^{\pi^{k}}_{h}(s,a)}{\tilde{\mu}^{*}_{h}(s,a)}\leq\frac{5\hat{d}^{\pi^{k}}_{h}(s,a)}{3\hat{\mu}^{*}_{h}(s,a)}~{}~{}\forall k\in[K]$.
Proof.
First,$|\hat{\mu}_{h}^{*}(s,a)-\tilde{\mu}_{h}^{*}(s,a)|=|\sum_{k=1}^{K}\hat{\alpha}^{*}_{k}(\hat{d}^{\pi^{k}}_{h}(s,a)-d^{\pi^{k}}_{h}(s,a))|\leq\sum_{k=1}^{K}\hat{\alpha}^{*}_{k}|\hat{d}^{\pi^{k}}_{h}(s,a)-d^{\pi^{k}}_{h}(s,a)|\leq\frac{1}{4}\sum_{k=1}^{K}\hat{\alpha}^{*}_{k}d^{\pi^{k}}_{h}(s,a)=\frac{\tilde{\mu}^{*}_{h}(s,a)}{4}$. Second, $\frac{d^{\pi^{k}}_{h}(s,a)}{\tilde{\mu}^{*}_{h}(s,a)}\leq\frac{4\hat{d}^{\pi^{k}}_{h}(s,a)}{3\tilde{\mu}^{*}_{h}(s,a)}\leq\frac{5\hat{d}^{\pi^{k}}_{h}(s,a)}{3\hat{\mu}^{*}_{h}(s,a)}$. ∎
4.3 Estimation of the importance weighting ratio
@@
In this section, we introduce our algorithm for estimating the importance weighting ratios which is sketched out in Algorithm 1. Our algorithm is based on the idea of DualDICE Nachum et al. (2019). In DualDICE, they propose the following loss function
$\ell^{\pi}(w)=\frac{1}{2}\mathbb{E}_{s,a\sim\mu}\left[w^{2}(s,a)\right]-\mathbb{E}_{s,a\sim d^{\pi}}\left[w(s,a)\right],$ (7)
@@
the optimal minimum is achieved at $w^{\pi,*}(s,a)=\frac{d^{\pi}(s,a)}{\mu(s,a)}$ which is the distribution ratio. They address the on-policy limitation of the second term in 7 by changing the variable based on Bellman’s equation. However, their method only works for infinite horizon MDPs and it becomes unclear how to optimize the loss function after the variable change. We propose a new step-wise loss function which works for finite horizon MDPs. More importantly, the loss function is strongly-convex and smooth, hence, we can give finite sample results based on the convergence rate in the optimization literature.
@@
Specifically, we define a loss function at each step and optimize them from step $h=1$ to step $h=H$. We define the loss function of policy $\pi$ at step $h$ as,
$\ell^{\pi}_{h}(w)$ $=\frac{1}{2}\sum_{s,a}\frac{\tilde{\mu}_{h}(s,a)}{\hat{\mu}_{h}(s,a)}w^{2}(s,a)-\sum_{s,a}\sum_{s^{\prime},a^{\prime}}\tilde{\mu}_{h-1}(s^{\prime},a^{\prime})P(s|s^{\prime},a^{\prime})\pi(a|s)\frac{\hat{w}_{h-1}(s^{\prime},a^{\prime})}{\hat{\mu}_{h-1}(s^{\prime},a^{\prime})}w(s,a)$
$=\frac{1}{2}\mathbb{E}_{s,a\sim\tilde{\mu}_{h}}\left[\frac{w^{2}(s,a)}{\hat{\mu}_{h}(s,a)}\right]-\mathbb{E}_{s^{\prime},a^{\prime}\sim\tilde{\mu}_{h-1},s\sim P_{h-1}(\cdot|s^{\prime},a^{\prime})}\left[\sum_{a}\frac{\hat{w}_{h-1}(s^{\prime},a^{\prime})}{\hat{\mu}_{h-1}(s^{\prime},a^{\prime})}w(s,a)\pi(a|s)\right]$
@@
where $\tilde{\mu}_{h}=\sum_{k=1}^{K}\hat{\alpha}_{k}d^{\pi^{k}}_{h}$ is the sampling distribution, and $\hat{\mu}_{h}=\sum_{k=1}^{K}\hat{\alpha}_{k}\hat{d}^{\pi^{k}}_{h}$ is the optimal solution of the approximate opt problem (refer to Section 4.2) and we set $\tilde{\mu}_{0}$ as empty, $\hat{w}_{0}=\hat{\mu}_{0}=1$ for notation simplicity.
@@
It is clear that the loss function is strongly convex and smooth. An even better property about our loss function is that the smoothness factor $\xi$ and strongly-convexity factor $\gamma$ are well bounded based on the property of our coarse distribution estimator, i.e. $\frac{4}{5}\leq\frac{\tilde{\mu}_{h}(s,a)}{\hat{\mu}_{h}(s,a)}\leq\frac{4}{3}$ which is a trivial conclusion from Lemma 4.4,
$\xi=\max_{s,a}\frac{\tilde{\mu}_{h}(s,a)}{\hat{\mu}_{h}(s,a)},\ \gamma=\min_{s,a}\frac{\tilde{\mu}_{h}(s,a)}{\hat{\mu}_{h}(s,a)},\ \frac{\xi}{\gamma}\leq\frac{5}{3}.$ (8)
This property actually plays an important role in the final sample complexity which we will revisit later. In the following lemma, we show that our step-wise loss function has an another nice property on step-to-step error propagation.
@@
Proof.
See Appendix A.4. ∎
Before introducing the sample complexity of estimating the distribution ratios by Algorithm 1, we first give the following convergence rate of minimizing strongly-convex loss functions with stochastic gradient descent.
Algorithm 1 Importance Density Estimation ($\mathrm{IDES}$ )
  Input: Horizon $H$, feasible sets $\{D_{h}\}_{h=1}^{H}$, accuracy $\epsilon$, target policy $\pi$, coarse estimator $\{\hat{d}^{\pi}_{h}\}_{h=1}^{H}$ and $\{\hat{\mu}_{h}\}_{h=1}^{H}$.
  Initialize $w_{h}^{0}=0,\ h=1,\dots,H$ and assume $\mu_{0}=Empty$ for simple presentation.
  for $h=1$ to $H$ do
     for $i=1$ to $n_{h}$ do
        Calculate gradient $g(w_{h}^{i-1})$,
$g(w_{h}^{i-1})(s,a)=\frac{w_{h}^{i-1}(s,a)}{\hat{\mu}_{h}(s,a)}\mathbb{I}(s^{i}_{h}=s,a^{i}_{h}=a)-\frac{\hat{w}_{h-1}(s^{i}_{h-1},a^{i}_{h-1})}{\hat{\mu}_{h-1}(s^{i}_{h-1},a^{i}_{h-1})}\pi(a|s)\mathbb{I}(s^{i^{\prime}}_{h}=s)$
        Update $w_{h}^{i}=Proj_{w\in D_{h}}\{w_{h}^{i-1}-\eta_{h}^{i}g(w_{h}^{i-1})\}$.
     end for
  end for
Lemma 4.6 (Theorem 25.2 in Notes).
$\mathbb{E}[L(\hat{w})-L(w^{*})]\leq\frac{2G^{2}}{\lambda(T+1)},$
where $G^{2}$ is the variance bound of the stochastic gradient.
Based on the above results, implementing Algorithm 1 with iterations $n_{h}$ specified in the algorithm, we can achieve an accurate estimation of the distribution ratios which we formalize in the following lemma.
@@
Implement Algorithm 1 with parameters $n_{h}$ specified in the algorithm, we have,
$\mathbb{E}\left[\sum_{s,a}\left|\tilde{\mu}_{h}(s,a)\frac{\hat{w}_{h}(s,a)}{\hat{\mu}_{h}(s,a)}-d^{\pi^{k}}_{h}(s,a)\right|\right]\leq\frac{\epsilon}{4H}.$
Proof.
See Appendix A.5 for complete proof. Here we provide proof sketch. By Lemma 4.5 and the property of smoothness $\|\nabla\ell^{\pi}_{h}(\hat{w}_{h})\|_{1}^{2}\leq 2\xi(\ell^{\pi}_{h}(\hat{w}_{h})-\ell^{\pi}_{h}(w^{*}_{h}))$, we need to minimize the loss function at each step $h$ such that $\ell^{\pi}_{h}(\hat{w}_{h})-\ell^{\pi}_{h}(w^{*}_{h})\leq\frac{\epsilon^{2}}{32\xi H^{4}}$. Invoking the convergence rate in Lemma 4.6, we need samples $n=O\left(\frac{\xi}{\gamma}\frac{H^{4}G^{2}}{\epsilon^{2}}\right)$. Remember that we have shown in (8) that $\frac{\xi}{\gamma}\leq\frac{5}{3}$, this nice property helps us to get rid of the undesired ratio of the smoothness factor and the strongly-convexity factor, i.e. $\frac{\max_{s,a}\mu(s,a)}{\min_{s,a}\mu(s,a)}$ of the original loss function (7) which can be extremely bad. Finally, replacing $G^{2}$ by $O\left(\sum_{s,a}\frac{(\hat{d}_{h}^{\pi}(s,a))^{2}}{\hat{\mu}_{h}(s,a)}+\frac{(\hat{d}_{h-1}^{\pi}(s,a))^{2}}{\hat{\mu}_{h-1}(s,a)}\right)$ which is the variance bound of the stochastic gradient in our case. ∎

### v2

@@ 3.1 Contributions
Our main contribution is to propose an algorithm named $\mathrm{CAESAR}$ for multiple-policy evaluation using two phases. In the first phase, we coarsely estimate the visitation distributions of all deterministic policies at the cost of a lower-order sample complexity. In the second phase, with the coarse distribution estimators, we can solve a convex optimization problem to build an approximately optimal sampling distribution $\tilde{\mu}^{*}$ with which we estimate the performance of target policies using marginal importance weighting. Using $\mathrm{CAESAR}$ , a number of trajectories $n=\tilde{O}\left(\frac{H^{4}}{\epsilon^{2}}\sum_{h=1}^{H}\max_{k\in[K]}\sum_{s,a}\frac{(d_{h}^{\pi^{k}}(s,a))^{2}}{\mu^{*}_{h}(s,a)}\right)$, and probability at least $1-\delta$, we can evaluate the performance of all target policies up to an $\epsilon$ error. $\mathrm{CAESAR}$ is consistently better than the naive uniform sampling strategy over target policies as described in (2). $\mathrm{CAESAR}$ also improves upon the result (1) by Dann et al. [5] in some cases where their results have a dependency on $K$ while ours do not (see Section 5).
In addition to our main contribution, we propose two sub-algorithms that are of independent interest. First, we propose $\mathrm{MARCH}$ which achieves coarse estimation of all deterministic policies with sample complexity $\tilde{O}(\frac{\mathrm{poly}(H,S,A)}{\epsilon})$, even though the number of all deterministic policies is exponential. Second, we propose $\mathrm{IDES}$ to accurately estimate the marginal importance ratio by minimizing a carefully designed step-wise loss function using stochastic gradient descent, which is modified from the idea of DualDICE [21]. Besides, we also utilize a Median-of-Means estimator [20] to convert the in-expectation result to a high-probability result.
@@ 4 Main Results and Algorithm
In this section, we introduce $\mathrm{CAESAR}$ which is sketched out in Algorithm 2 and present the main results. Different from on-policy evaluation, we build a single sampling distribution with which we can estimate the performance of all target policies using importance weighting. To that end, we first coarsely estimate the visitation distributions of all deterministic policies at the cost of a lower-order sample complexity. Based on these coarse distribution estimators, we can build an optimal sampling distribution by solving a convex optimization problem. Finally, we utilize the idea of DualDICE [21] with some modifications to estimate the importance-weighting ratio. In the following sections, we explain the steps of $\mathrm{CAESAR}$ in detail.
4.1 Coarse estimation of visitation distributions
We first introduce a proposition that shows how we can coarsely estimate the visitation distributions of target policies with lower-order sample complexity $\tilde{O}(\frac{1}{\epsilon})$. Although this estimator is coarse and cannot be used to directly evaluate the performance of policies, which is our ultimate goal, it possesses nice properties that enable us to construct the optimal sampling distribution and estimate the importance weighting ratio in the following sections.
The idea behind the feasibility of computing these estimators is based on the following lemma which shows that estimating the mean value of a Bernoulli random variable up to constant multiplicative accuracy only requires $\tilde{O}(\frac{1}{\epsilon})$ samples.
@@ 4.1 Coarse estimation of visitation distributions
Let $Z_{\ell}$ be i.i.d. samples $Z_{\ell}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathrm{Ber}(p)$. Setting $t\geq\frac{C\log(C/\epsilon\delta)}{\epsilon}$, for some known constant $C>0$, it follows that with probability at least $1-\delta$, the empirical mean estimator $\hat{p}_{t}=\frac{1}{t}\sum_{\ell=1}^{t}Z_{\ell}$ satisfies, $|\hat{p}_{t}-p|\leq\max\{\epsilon,\frac{p}{4}\}$.
@@
Proposition 4.2 works by running each policy independently and applying Lemma 4.1. However, this would induce an exponential dependency on $S,A$, if for example we aim to coarsely estimate all deterministic policies. To fix this we propose an algorithm named $\mathrm{MARCH}$ (see Appendix B) that leverages the overlapping visitations of the policy set. Through a novel analysis, we show that $\mathrm{MARCH}$ achieves coarse estimation of all deterministic policies with sample complexity $\tilde{O}(\frac{\mathrm{poly}(H,S,A)}{\epsilon})$.
@@
Based on Lemma 4.3, we can ignore the state-action pairs satisfying $\hat{d}(s,a)\leq 5\epsilon^{\prime}$. Since if we replace $\epsilon^{\prime}$ by $\frac{\epsilon}{14SA}$, the error of performance estimation induced by ignoring these state-action pair is at most $\frac{\epsilon}{2}$. For simplicity of presentation, we can set $\hat{d}^{\pi}(s,a)=d^{\pi}(s,a)=0$ if $\hat{d}^{\pi}(s,a)<\frac{5\epsilon}{14SA}$. Hence, we have that,
@@ 4.2 Optimal sampling distribution
We evaluate the expected total rewards of target policies by importance weighting, using samples $\{s_{1}^{i},a_{1}^{i},s_{2}^{i},a_{2}^{i},\dots,s^{i}_{H},a^{i}_{H}\}_{i=1}^{n}$ drawn from a sampling distribution $\{\mu_{h}\}_{h=1}^{H}$. Specifically, $\hat{V}^{\pi^{k}}_{1}=\frac{1}{n}\sum_{i=1}^{n}\sum_{h=1}^{H}\frac{d_{h}^{\pi^{k}}(s_{h}^{i},a_{h}^{i})}{\mu_{h}(s_{h}^{i},a_{h}^{i})}r_{h}(s^{i}_{h},a^{i}_{h}),\ k\in[K]$. To minimize the variance of our estimator (see Appendix A.2), we find the optimal sampling distribution by solving the following convex optimization problem,
@@
However, in some cases, the optimal $\mu^{*}$ may not be realized by any policy (see Appendix A.3). Therefore, to facilitate the construction of the sampling distribution $\mu^{*}$, we constrain $\mu_{h}$ to lie within the convex hull of $\mathcal{D}=\{d^{\pi}_{h}:\pi\in\Pi_{det}\}$ which formulates the constrained optimization problem,
$\mu^{*}_{h}=\operatorname*{arg\,min}_{\mu\in\text{conv}(\mathcal{D})}\max_{k\in[K]}\sum_{s,a}\frac{(d_{h}^{\pi^{k}}(s,a))^{2}}{\mu(s,a)},\ h\in[H].$ (5)
We denote the optimal solution to (5) as $\mu^{*}_{h}=\sum_{\pi\in\Pi_{det}}\alpha^{*}_{\pi}d^{\pi}_{h}$. Since $d^{\pi^{k}}_{h}$ is unknown, we can only solve the approximate optimization problem,
$\hat{\mu}^{*}_{h}=\operatorname*{arg\,min}_{\mu\in\text{conv}(\mathcal{\hat{D}})}\max_{k\in[K]}\sum_{s,a}\frac{(\hat{d}_{h}^{\pi^{k}}(s,a))^{2}}{\mu(s,a)},\ h\in[H],$ (6)
where $\mathcal{\hat{D}}=\{\hat{d}^{\pi}_{h}:\pi\in\Pi_{det}\}$. We denote the optimal solution to (6) by $\hat{\mu}^{*}_{h}=\sum_{\pi\in\Pi_{det}}\hat{\alpha}^{*}_{\pi}\hat{d}^{\pi}_{h}$. Correspondingly, our real sampling distribution would be $\tilde{\mu}^{*}_{h}=\sum_{\pi\in\Pi_{det}}\hat{\alpha}^{*}_{\pi}d^{\pi}_{h}$.
@@
If property (3) holds: $|\hat{d}_{h}^{\pi^{k}}(s,a)-d^{\pi^{k}}_{h}(s,a)|\leq\frac{d_{h}^{\pi^{k}}(s,a)}{4},\ \forall s\in\mathcal{S},a\in\mathcal{A},h\in[H],k\in[K]$, then $|\tilde{\mu}_{h}^{*}(s,a)-\hat{\mu}_{h}^{*}(s,a)|\leq\frac{\tilde{\mu}_{h}^{*}(s,a)}{4}$.
4.3 Estimation of the importance density
Algorithm 1 Importance Density Estimation ($\mathrm{IDES}$ )
  Input: Horizon $H$, accuracy $\epsilon$, target policy $\pi$, coarse estimator $\{\hat{d}^{\pi}_{h}\}_{h=1}^{H}$ , $\{\hat{\mu}_{h}\}_{h=1}^{H}$ and feasible sets $\{D_{h}\}_{h=1}^{H}$ where $D_{h}(s,a)=[0,2\hat{d}^{\pi}_{h}(s,a)]$.
  Initialize $w_{h}^{0}=0,\ h=1,\dots,H$, and set $\mu_{0}(s_{0},a_{0})=1,P_{0}(s|s_{0},a_{0})=\nu(s),\hat{w}_{0}=\hat{\mu}_{0}=1$.
  for $h=1$ to $H$ do
     for $i=1$ to $n_{h}$ do
        Calculate gradient $g(w_{h}^{i-1})$,
$g(w_{h}^{i-1})(s,a)=\frac{w_{h}^{i-1}(s,a)}{\hat{\mu}_{h}(s,a)}\mathbb{I}(s^{i}_{h}=s,a^{i}_{h}=a)-\frac{\hat{w}_{h-1}(s^{i}_{h-1},a^{i}_{h-1})}{\hat{\mu}_{h-1}(s^{i}_{h-1},a^{i}_{h-1})}\pi(a|s)\mathbb{I}(s^{i^{\prime}}_{h}=s).$
        Update $w_{h}^{i}=Proj_{w\in D_{h}}\{w_{h}^{i-1}-\eta_{h}^{i}g(w_{h}^{i-1})\}$.
     end for
  end for
@@ 4.3 Estimation of the importance density
In this section, we introduce the $\mathrm{IDES}$ algorithm for estimating the importance weighting ratios which is sketched out in Algorithm 1. $\mathrm{IDES}$ is based on the idea of DualDICE [21]. In DualDICE, they propose the following loss function
$\ell^{\pi}(w)=\frac{1}{2}\mathbb{E}_{s,a\sim\mu}\left[w^{2}(s,a)\right]-\mathbb{E}_{s,a\sim d^{\pi}}\left[w(s,a)\right].$ (7)
@@
The minimum of $\ell^{\pi}(\cdot)$ is achieved at $w^{\pi,*}(s,a)=\frac{d^{\pi}(s,a)}{\mu(s,a)}$, the importance weighting ratio. In [21] they tackle the on-policy limitation of the second term in (7) by transforming the variable based on Bellman’s equation. However, their method only works for infinite horizon MDPs and it becomes unclear how to optimize the loss function after the variable change. We propose a new step-wise loss function which works for finite horizon MDPs. More importantly, the loss function is strongly-convex and smooth, enabling optimization through stochastic gradient descent and yielding non-asymptotic sample complexity results.
@@
More precisely, we define the step-wise loss function of policy $\pi$ at each step $h$ as,
$\ell^{\pi}_{h}(w)=\frac{1}{2}\mathbb{E}_{s,a\sim\tilde{\mu}_{h}}\left[\frac{w^{2}(s,a)}{\hat{\mu}_{h}(s,a)}\right]-\mathbb{E}_{s^{\prime},a^{\prime}\sim\tilde{\mu}_{h-1},s\sim P_{h-1}(\cdot|s^{\prime},a^{\prime})}\left[\sum_{a}\frac{\hat{w}_{h-1}(s^{\prime},a^{\prime})}{\hat{\mu}_{h-1}(s^{\prime},a^{\prime})}w(s,a)\pi(a|s)\right]$
@@
where $\tilde{\mu}_{h}=\sum_{\pi\in{\Pi_{det}}}\hat{\alpha}^{*}_{\pi}d^{\pi}_{h}$ is the sampling distribution, and $\hat{\mu}_{h}=\sum_{\pi\in\Pi_{det}}\hat{\alpha}^{*}_{\pi}\hat{d}^{\pi}_{h}$ is the optimal solution to the approximate optimization problem (6), and we set $\tilde{\mu}_{0}(s_{0},a_{0})=1,P_{0}(s|s_{0},a_{0})=\nu(s),\hat{w}_{0}=\hat{\mu}_{0}=1$ for notational simplicity.
@@
This loss function possesses two nice properties. First, it is $\gamma$-strongly convex and $\xi$-smooth where $\gamma=\min_{s,a}\frac{\tilde{\mu}_{h}(s,a)}{\hat{\mu}_{h}(s,a)},\xi=\max_{s,a}\frac{\tilde{\mu}_{h}(s,a)}{\hat{\mu}_{h}(s,a)}$. Based on the property of our coarse distribution estimator, i.e., $\frac{4}{5}\leq\frac{\tilde{\mu}_{h}(s,a)}{\hat{\mu}_{h}(s,a)}\leq\frac{4}{3}$, which is a trivial corollary from Lemma 4.4, $\gamma$ and $\xi$ are bounded as well as their ratio, i.e. $\frac{\xi}{\gamma}\leq\frac{5}{3}$. This property actually plays an important role in deriving the final sample complexity, which we discuss in Appendix A.5.
In the following lemma, we show that our step-wise loss function has friendly step-to-step error propagation properties.
@@
Lemma 4.5 indicates that using the distribution ratio estimator from the previous step allows us to estimate the ratio at the current step, introducing only an additive error. Consequently, by optimizing step-by-step, we can achieve an accurate estimation of the distribution ratios at all steps, as formalized in the following lemma.
@@
Implementing Algorithm 1, we have the importance density estimator $\frac{\hat{w}_{h}(s,a)}{\hat{\mu}_{h}(s,a)}$ such that,
$\mathbb{E}\left[\sum_{s,a}\left|\tilde{\mu}_{h}(s,a)\frac{\hat{w}_{h}(s,a)}{\hat{\mu}_{h}(s,a)}-d^{\pi^{k}}_{h}(s,a)\right|\right]\leq\frac{\epsilon}{4H},\ h\in[H].$ (8)